# PHYSICISTS JOKE

### v1

@@ 3.5 Reports I Have Read… And Perhaps Written
The technical report as a specialized form of scientific literature has risen from obscurity in recent years to occupy an exceptionally important position in the field of information exchange. Ideally, the main task of every report is to transmit scientific information—to transmit it accurately, neatly, and unambiguously. Alas, in practice, all "reports" break down into a multitude of categories based on their degree of approximation to this ideal, as well as to specific reasons that prevent them from achieving it. Below are described and classified the most important categories with which the author has had to deal during his fifteen years of activity in the field of scientific information. Regarding the most senseless reports, it should be said that they do not break down into clear, mutually exclusive groups. I chose the names based on the dominating feature, but on the whole, such reports are "not a proud people" and easily adopt bad habits from one another.
@@ 1. The Mystery Report
When reading such a work, it seems that the author almost intentionally tries to keep the reader in ignorance regarding what, exactly, it is written about as long as possible. In some cases, the success is complete, and the riddle remains unsolved to the very end. The mystery, of course, thickens even more if there is an absolutely uninformative title on the document. I saw two such titles recently while leafing through an abstract journal: "Review of Work" and "Current Work. Report XXIX".
Mystery reports usually begin with some historical statement like: "In 1927, Professor K. K. McJillicuddy, working at ABC University, discovered such-and-such." This initial phrase sounds approximately the same as "In a faraway kingdom, in a certain state…", with which all children’s fairy tales begin, and it requires about the same mental effort from its inventor. A similar beginning is, of course, an excellent exit for the author of the report if he is not very sure about what he wants to say, or if he is simply too lazy to start dirtying paper. For the reader, such a beginning introduces an element of mystery from the very first sentence. Beyond the initial assertion, further digressions on a historical theme may follow, and behind them other materials, only indirectly related to the true purpose of this document, which hangs in the air. And the reader plunges into the labyrinth of the report, having not the slightest idea of whether the arrows leading to the exit door actually exist, and not being even sure that this door really exists. As a rule, the reader eventually digs down to what the author was trying to tell him. However, this feat usually requires such a tension of analytical abilities and such mental subtlety that it no longer delivers joy to the completely exhausted reader.
2. The "Say It Again, Please" Report
At the disposal of anyone who wishes to learn how to write reports of this type, there is a wide selection of effective methods. First of all, the author, of course, must beware of simple non-propagated sentences. The unholy trio of subject, predicate, and object is too direct a means of transmitting information, and therefore does not satisfy. Instead of this, the author everywhere, where only possible, must use long, ornate, flowery phrases, diluting them with introductory words and starting with subordinate clauses. I will allow myself to cite as an illustration a short excerpt from a dialogue that took place several years ago at a meeting of one of the Congressional committees. An official of the Ministry of Defense was asked if their department plans the construction of a certain underground structure. Here is what he answered:
@@ 2. The "Say It Again, Please" Report
"We are trying to maintain a balance between stationary, as they are sometimes called, installations, the creation of which in individual cases may be associated with the construction of underground structures, on the one hand, and the effectiveness of our defensive means, on the other hand, which, obviously, according to the views adopted at the present time, is considered as the strongest argument against the construction of closed defensive installations."
After which one of the committee members stated: "You are simply magnificent. I didn’t understand a damn thing in this verbal eruption. For God’s sake, what did you mean?"
This example is good not only because it contains a brilliant specimen of the style "Repeat, please, one more time." The reaction of the "consumer" to such semantic twists is also completely typical.
Not wishing to create the impression that only government officials possess a monopoly on such sayings, I will tell about a less official case. The quotation I cite below is borrowed from the article "Ascending and Descending Air Currents."
(First sentence) "While the speed of the flow remains at an average level—lower than in the adiabatic case for dry air, and higher than in the adiabatic case for air saturated with water vapor, we can easily imagine that an isolated air mass, which is saturated to absolute humidity at a temperature slightly exceeding the temperature of the surrounding air, will turn out to be in a state to begin an ascending movement, since under the postulated conditions its temperature at any achieved height will be higher than the temperature of the surrounding air."
(Second sentence) "However, the opposite process—a descending flow—is not so easy to understand."
To the credit of the famous encyclopedia from which this quote is taken, it should be noted that in the next, revised edition, the second sentence was redone.
To achieve the worst results, the author must strive to maintain two numerical coefficients at an extremely high level—the "number of syllables per word" and the "average number of words per sentence." In the first case, follow the motto: "Never use a monosyllabic word if there is a synonym of six or seven syllables." Your report will especially win if, in addition, you use these long words out of place.
There are many ways to maintain the second index at a high level. For example, "bearing in mind the fact that" is always better than simply "because," "equal in order of magnitude" is much more effective than "about." You are free to start at least every sentence with such common (and such meaningless) turns of phrase as "it should be noted that," "it is not uninteresting to pay attention to the fact that," etc. These verbiages give everything stated such an impersonal character, as if removing any responsibility from the author for their content. A person, conscientiously following these recipes and supplementing them with the ubiquitous use of the passive voice and professional jargon everywhere, where only possible, and flavoring the whole dish with non-germane definitions and additions, can be completely calm about the result.
@@ 3. The "Missing Thought" Report
When reading such reports, tests on ingenuity immediately come to mind, with which we were tormented in school years. Only instead of missing words, the reader must restore missing thoughts, that is, significant chunks of information and argumentation which the author did not consider necessary to include in the text. In such cases, the reader, trying to independently make that step which appeared obvious to the author, most often finds himself in the position of students at a lecture on mathematics, about which I recently read. The professor, standing at the blackboard, was immersed in a lengthy derivation. At some point, he pronounced the standard phrase "from here it clearly follows the following" and wrote a long and complex expression, absolutely unlike anything written earlier. Then he hesitated, a puzzled expression appeared on his face, he muttered something and went out of the auditorium into his office. Appearing from there half an hour later, he explained to the audience with a satisfied look: "I was right. It is, indeed, completely obvious."
Reading such a report, you, figuratively speaking, slide smoothly along a smooth rail track following the train of the author’s thought and suddenly stumble upon a destroyed section or a precipice. You clearly see that the track continues on the other side of the abyss, but there is no bridge, and there is nothing to build it from.
4. Camouflaging Reports
Characteristic features (they can occur all at once or in certain combinations):
@@ 4. Camouflaging Reports
    The results are not presented fully.
@@
    The conclusions obtained do not follow from the results.
@@
    The recommendations cited do not follow from the conclusions.
@@
In finished, unmistakable samples of camouflaging reports, the author, pretending to be a scientist, acts in reality like a hawker. He wants to sell something in the literal sense or seduce the buyer with some idea of his. Camouflaging reports of a high class can be recognized immediately by their beautiful binding, often even with gold lettering, by the magnificent illustrations not less than in four colors, interleaved with tissue paper, and by the glossy expensive paper.
On this, perhaps, one can finish the discussion of typical "freaks" in the family of technical reports, the main task of which, as I said in the introduction and repeat now, is the transmission of information—transmission that is accurate, neat, and unambiguous.
@@ 4.1 How to Conduct Scientific Research
Much research has been devoted to the art of placing one’s name at the head of the list of authors, but some subtle questions remain unilluminated. The Alphabetical Trick. Since alphabetical order in compiling lists of authors is gradually becoming generally accepted, it is useful to concentrate on creating a solid advantage for oneself personally. This can be achieved in two ways: change your surname so that the new one begins with the letter A, or choose co-authors with surnames from the lower half of the alphabet (see, for example, G. C. Wick, A. S. Wightman, E. C. Wigner, Phys. Rev., 88, 101, 1952). But in this case, it is easy to miss the mark. One should not chase co-authors of too large a caliber. An article by A. Bar-Nudnik and Albert Einstein will always be referred to as "Einstein et al.", regardless of the order of names.
@@ 4.2 The Computer Poet
4.3 Preparation of Coffee in Research Institutions
@@
A significant part of the time in laboratories and offices is devoted by scientists to the preparation and consumption of coffee (in Great Britain they prefer tea). The amount of beverage consumed and the time spent on consumption range from one cup and five minutes a day to several dozen cups and several hours. There is a very acute need to regulate this activity. This note represents the first attempt to generalize the rich experience currently going to waste.
@@
All commercially available varieties of coffee are used (with the exception of acorn and synthetic), of fine, medium, normal, and coarse grind, as well as in beans. Boiling is performed in vessels made of glass, aluminum, or stainless steel, including laboratory beakers, fireproof flasks, strainers, autoclaves, and (there was once a case) distillation stills. However, ordinary coffee pots are also used. Sources of energy for raising the temperature of the extracting water can be the flame of a gas burner, electric current, superheated steam, exhaust gases of an internal combustion engine, and the reaction of alcohol oxidation (do not use ether – it is explosive). In summer, cooling systems are used: refrigerators, freezers, ice cubes, dry ice, and liquid air.
Methods
@@
    "Amateur". Coarse-ground coffee is poured into cold raw water, then the water is brought to a boil. The sediment is thoroughly shaken, and the resulting suspension is poured into cups. That the resulting product is indeed coffee can perhaps only be verified with the help of photometric measurements. The same applies to a slightly more economical modification of the amateur method, where the sediment is filtered off and saved for reuse.
@@
    "Professional". Water is heated to $99^{\circ}$C, fine-ground coffee is added (approximately 1.2 grams per cup), the liquid is brought to a boil and removed from the fire. Centrifuging the sediment is not necessary. The resulting drink has a delicate taste, which sometimes can be killed by adding milk.
@@
    "Expert". Statistical processing of tasting results allowed for an important improvement: water is first brought to a boil, then the flame is reduced to a minimum, and coffee is poured into the water (one spoon per cup). Everything is boiled for 10 seconds, then settles for a few minutes, and one can drink.
@@
    "Express" (should be distinguished from the "Expresso" method described below). Instant coffee is used in combination with boiling or hot water according to the instructions written on the tin. This method has two advantages over all others: speed and the absence in this case of the notorious "wonderful aroma of fresh coffee," which allows avoiding an invasion of the thirsty from the neighboring laboratory.
@@
    "Expresso". This name has already become a household word for steam extraction, when superheated steam is passed through compressed coffee powder and then cooled. The resulting condensate possesses the color and smell of coffee. In laboratory conditions, one can use the "Soxhlet" extraction apparatus, however, this method is too labor-intensive for use, except in cases where you have a surplus of technical personnel, a lack of ideas, and you cannot think of a more reasonable way to use the apparatus.
@@ 4.5 Advice to the Examiner
Original Source: Electronics.
@@
    First of all, explain to the examinee that his entire professional career may collapse because of his unsuccessful answer. Emphasize the importance of the situation to him. Put him in his place from the very beginning.
@@
    Ask the hardest questions right away. If the first question is sufficiently difficult or confused, the examinee will get too nervous to answer the following questions, no matter how simple they are.
@@
    Addressing the examinee, maintain restraint and dryness; with other examiners, however, be very cheerful. It is effective to address other examiners from time to time with mocking remarks regarding the examinee’s answers, ignoring him himself, as if he is not in the room.
@@
    Force the examinee to solve problems by your method, especially if this method is unusual. Limit the examinee by inserting a multitude of instructions and reservations into every question. The idea here consists of complicating a task which without this would be quite simple.
@@
    Force the examinee to make a trivial error, and let him rack his brains over it as long as possible. Immediately after he notices the mistake, but just before he understands how to fix it, contemptuously correct him yourself. This requires high insight and precise timing, which is achieved only by great practice.
@@
    When the examinee begins to drown, never help him scramble out. Yawn… and move on to the next question.
@@
    Ask the examinee questions from time to time like: "Didn’t you cover this in elementary school?"
@@
    Do not allow the examinee to ask clarifying questions and never repeat your own explanations and statements.
@@
    Every few minutes ask if he is worried.
@@
    Wear dark glasses. Impenetrability makes people nervous.
@@
    Finishing the exam, say to the examinee: "Wait outside the door. We will call you."

### v2

@@ 3.5 Reports I Have Read… And Perhaps Written
The technical report, as a specialized form of scientific literature, has risen from relative obscurity in recent years to a position of immense importance in the dissemination of information. Ideally, the function of every report is to convey information—to convey it accurately, clearly, and unambiguously. Alas, in practice, "reports" fall into a variety of categories based on their degree of approximation to this ideal. described below are a few of the more distinctive species I have encountered during fifteen years in the information business.
@@ 1. The Mystery Report
The author of this document seems to be trying to keep the reader in suspense as long as possible regarding what the report is actually about. In extreme cases, he succeeds in maintaining the secret to the bitter end. The mystery is, of course, deepened if the report bears a non-committal title like "Progress Report" or "Summary of Work."
Mystery Reports usually begin with historical statements such as: "In 1927, Professor K. K. McJillicuddy, working at ABC University, discovered such-and-such." This opening strikes a note reminiscent of the "Once upon a time…" of children’s fairy tales. For the author, it is an excellent device if he is not quite sure what he wants to say, or is simply reluctant to start saying it. For the reader, it establishes an atmosphere of uncertainty immediately. The reader plunges into the body of the report not knowing whether the arrows lead to an exit, or indeed if there is an exit. Usually, the reader eventually unearths the author’s message. However, the feat typically requires such a strain on his deductive powers that he has little energy left for rejoicing.
2. The "Say It Again" Report
The aspirant to this category has a wide choice of methods. First, he must avoid simple, declarative sentences. The unholy trio of subject, predicate, and object is far too direct; instead, the author should use long, involved, circuitous phrases, qualifying them with parenthetical insertions and starting with dependent clauses. To illustrate, I quote a sentence from a dialogue that occurred some years ago in a Congressional committee hearing. A Defense Department official was asked if his agency planned to build a certain underground structure. He replied:
@@ 2. The "Say It Again" Report
"We are attempting to maintain a balance between static, as they are sometimes called, installations, the creation of which in individual cases may be associated with the construction of underground structures, on the one hand, and the effectiveness of our defensive means, on the other hand, which, obviously, according to views adopted at the present time, is considered as the strongest argument against the construction of closed defensive installations."
To which a Congressman replied: "You are magnificent! I didn’t understand a word of that eruption. For God’s sake, what do you mean?"
To achieve the worst results, the author must strive to keep two indices high: the "syllable-per-word" count and the "words-per-sentence" count. The motto is: "Never use a monosyllable if a synonym of six or seven syllables is available." Sentences should begin with such meaningful phrases as "It is interesting to note that…" or "Attention should be called to the fact that…" These verbal paddings give the report an impersonal flavor, as if absolving the author of responsibility for the contents.
@@ 3. The "Missing Thought" Report
Reading this type of report reminds one of the "fill-in-the-blank" tests we took in school. Only here, instead of missing words, the reader must supply missing logic—significant chunks of information and argumentation which the author did not feel necessary to include. The reader finds himself in the position of the students in the famous mathematics lecture story. The professor, covering the blackboard with a derivation, arrives at a point where he says, "From this, it is obvious that…" and writes a long, complex expression bearing no resemblance to what came before. He pauses, looks puzzled, mutters, and walks into his office. Thirty minutes later he returns, beaming: "I was right. It is obvious."
In reading such a report, you glide smoothly along the rails of the author’s logic only to find a washed-out bridge. You can see the track continuing on the other side of the chasm, but there is no bridge, and no lumber to build one.
4. The Camouflage Report
Characteristic features (occurring singly or in combination):
@@ 4. The Camouflage Report
    Data are presented, but not analyzed.
@@
    Conclusions are drawn, but do not follow from the data.
@@
    Recommendations are made, but do not follow from the conclusions.
@@
In the most finished examples of the Camouflage Report, the author, ostensibly a scientist, is actually a pitchman. He wants to sell something. High-class Camouflage Reports can be recognized immediately by their expensive covers (often with gold lettering), magnificent four-color illustrations separated by tissue paper, and glossy, high-quality paper.
Here we may conclude our discussion of the "freaks" of the report family. The true function of a report remains, as stated in the introduction, the transmission of information—transmission that is accurate, clear, and unambiguous.
3.6 How to Use Slides
(Speech at the banquet marking the close of the International Conference on Nuclear Structure, Kingston, Canada, 1960) Author: D. G. Wilkinson
Original Source: Proceedings of the International Conference on Nuclear Structure (1960).
I have been asked to say a few words on an important question—how to use slides. It is difficult to initiate the amateur into all the subtleties of this art in a few minutes. Therefore, I intend to touch only upon the most elementary principles; one cannot expect more. I want to emphasize that my present communication is only an excerpt from the general "Rules of Conferencemanship" and is devoted to only one of the topics covered by that code. In such a short speech, I am deprived of the opportunity to touch upon such vital questions as: "How to mention your collaborators while implying that they do not deserve it," or "How to discredit the theory and experimental technique of your rival without understanding either one or the other."
The question of using slides breaks down into three sub-questions. On the third—"How to humiliate your opponent"—I was not permitted to speak. Two remain: "How to exhaust the projectionist" and "How to conquer the audience."
As regards the projectionist, the ultimate goal of the conferenceman is to drive him, if possible, to a nervous breakdown. It is important to establish the moment when you have succeeded in this, and then turn all attention to the listeners. The difficulty lies in the fact that, as a rule, you do not see the operator. But I believe it is usually sufficient to bring the process to the stage where his stuttering becomes audible in the hall; this has a useful unnerving effect on the audience. Such a state is self-sustaining, and after this, the operator can be left to himself.
Such primitive methods as using films of non-standard width or octagonal plates can be recommended only to the greenest novices. A satisfactory and more skilled opening is the "3–2–1" Technique. Here one exploits the fact that the operator always loads the first two slides into the apparatus while the chairman announces the talk, in order to switch on immediately. Instead of the expected "First slide, please," you say: "Third slide, please." (Elementary note: following this, you must quickly demand the second slide and only then—the first.)
The second method, best used in conjunction with the first, is the "White Spot Wobble." All slides are usually marked in one corner with a white spot, on which the operator must place his right thumb to ensure the correct orientation. You must place this spot in the wrong corner, so everything appears upside down. Applied together with the "3–2–1" technique, this acts stunningly. You show slight confusion, and then, brightening up, address the operator: "Oh, I beg your pardon, these plates are marked in an unusual way. You know, I usually take a personal operator with me." Then after some thought: "He is left-handed," and finally: "But don’t worry, only the first few shots are marked that way."
Immediately following this should be the "Parity-Violating Slide," which projects incorrectly no matter how you turn it in the frame. There are many ways to make such a slide. The simplest, yet exquisite: all letters should be drawn correctly, but words written from right to left.
Address the operator often. It is good if it is not quite clear whom you are actually addressing—him or the listeners. Avoid absurdly complicated instructions. Simply say: "In two slides I want to look again at the fourth from the end of those that have already been shown." And after the next frame add: "I, of course, meant that slide which will be fourth from the end of those shown after you show these two frames, and not the one which was fourth from the end when I spoke about it." After this, skip one slide.
These simple methods are sufficient for the majority of operators. In case of unexpected resistance, steeper measures can be taken. A grandmaster move is the "Thermal Jammer." To the bottom edge (on the screen, it will, of course, be the top) of a specially shortened slide, a thin, sufficiently elastic bimetallic strip is attached. When the image appears on the screen, you discuss it for some time and remark that the difference from the next curve is small, but will be striking if the slide is changed quickly enough. The bimetallic strip will have heated up by this time, bending the frame, and when you cry "Next, please!", it will inevitably jam firmly halfway. Driven by impatient requests of "Faster!", the operator will abandon attempts to move the frame by delicate tapping and will grab it with both hands, yanking it back and forth. The apparatus will crawl along the floor on all four legs, emitting pleasant screeching sounds. This always entertains the public.
Finally, the most sophisticated technique is the "Double Blank." Two absolutely clean slides are taken. They are placed after a series of frames which are demonstrated in rapid sequence, exerting a hypnotic effect. Suddenly this series ends, and the operator, having loaded the next pair, sighs with relief. However, the snapshot he projects after the next "Next, please," is one of the "Double Blanks," and the second frame in the holder is also empty. A few seconds later, an icy voice rings out: "I said: next, please!"—and the operator sees with horror that there is nothing on the screen, although he remembers perfectly well that he inserted the slide. Feeling his world collapsing, he pokes his finger straight into the middle of the clean slide to make sure it still exists. But a large hole has been made in the center of the slide beforehand; only the rim remains. Almost losing consciousness, the operator grabs the next frame from the box and tries to squeeze it into the holder, which is, naturally, occupied…
@@ 3.6 How to Use Slides
The Exam
At a physics exam, a professor writes an equation:
$E=h\nu$
and asks the student: "What is $\nu$?"
"Planck’s constant!"
"And hh?"
" The height of that plank!"
@@ 4.1 How to Conduct Scientific Research
Much research has been devoted to the art of placing one’s name at the head of the list of authors, but some subtle questions remain unilluminated. The Alphabetical Trick. Since alphabetical order in compiling lists of authors is gradually becoming generally accepted, it is useful to concentrate on creating a solid advantage for oneself personally. This can be achieved in two ways: change your surname so that the new one begins with the letter A, or choose co-authors with surnames from the lower half of the alphabet (see, for example, G. C. Wick, A. S. Wightman, E. C. Wigner, Phys. Rev., 88, 101, 1952). But in this case, it is easy to miss the mark. One should not chase co-authors of too large a caliber. An article by A. Bar-Nudnik and Albert Einstein will always be referred to as "Einstein et al.", regardless of the order of names. A more elegant solution to this problem is shown in Figure 4.1: arrange the authors in a circle, so that no one is first or last.
TITLEA. B. MishaymS. D. AdamE. F. OnyakhJ. G. BamadaQ. Z. QuackI. P. Freely
Figure 4.1: The “Circular Reference” method for determining authorship priority.
@@ 4.2 The Computer Poet
4.3 The Preparation of Coffee in Research Laboratories
@@
A considerable portion of the time in laboratories is devoted to the preparation and consumption of coffee. The following is a standardization of methods.
@@
All commercially available varieties of coffee are used (except acorn). Boiling is performed in borosilicate glass beakers, Erlenmeyer flasks, or, in emergencies, in distillation apparatus. Sources of heat include gas burners, electric hot plates, bunsen burners, and occasionally high-voltage arcs.
Procedures
@@
    The Crude Method. Coarse-ground coffee is poured into cold tap water, which is then brought to a boil. The sediment is suspended by vigorous shaking, and the resulting slurry is poured into cups. Verification that the product is indeed coffee can be achieved only by photometric analysis.
@@
    The Refined Method. Water is heated to $99^{\circ}$C. Fine-ground coffee is added (approx. 1.2 g per cup). The liquid is brought to a boil and removed from the flame. Centrifugation of the sediment is optional.
@@
    The "Expert" Method. Water is brought to a boil, the flame is reduced to minimum, and coffee is added. The mixture is boiled for exactly 10 seconds.
@@
    The "Espresso" Method. This term has become generic for steam extraction. In the laboratory, a Soxhlet extraction apparatus may be used, though this method is too labor-intensive unless one has a surplus of graduate students and a shortage of ideas.
@@
    The "Instant" Method. This involves the use of soluble powder. It has two advantages: speed, and the absence of the tell-tale aroma of fresh coffee, which prevents the invasion of thirsty colleagues from neighboring laboratories.
@@ 4.5 Advice to the Examiner
Original Source: Adapted from Electronics Magazine.
@@
    Explain to the candidate that his whole professional career depends on the outcome of this examination. Stress the gravity of the situation. Put him at his ease… near the edge of the chair.
@@
    Ask the hardest question first. If the first question is sufficiently obscure, the candidate will be too nervous to answer the subsequent simple ones.
@@
    Be reserved and stern with the candidate; be very cheerful with your fellow examiners. It is effective to address witty remarks to your colleagues about the candidate’s answers, ignoring him as if he were not in the room.
@@
    Force the candidate to solve problems by your method, especially if it is non-standard. Restrict him by inserting numerous conditions and reservations. The idea is to complicate a problem that would otherwise be quite simple.
@@
    Lead the candidate into a trivial error, and let him flounder as long as possible. Immediately after he notices the mistake, but just before he can correct it, correct him yourself with a look of disdain. This requires great timing.
@@
    When the candidate begins to drown, never help him out. Yawn… and move to the next question.
@@
    Ask questions like: “Didn’t you learn this in kindergarten?”
@@
    Never allow the candidate to ask clarifying questions, and never repeat your own explanations.
@@
    Every few minutes, ask him if he is nervous.
@@
    Wear dark glasses. Inscrutability is unnerving.
@@
    As the exam ends, say: “Wait outside. We’ll let you know.”